let {an} be a decreasing sequence of positive terms .prove that the series summation an sin (nX) converges uniformily on R if and only if nan tends to 0 as n tends to infinity
1
Expert's answer
2020-05-11T13:41:41-0400

Let "\{a_n\}" be the decreasing sequence of positive terms. Suppose "\sum_{n=1}^{\infty} a_n \sin(nx)"  converges uniformly and let the first "n" partial sum of the given series be "S_{n}(x)=\sum_{r=1}^{n} a_r \sin(rx)" , By the hypothesis "S_{n}(x)" converges uniformly.Now by Cauchy's criteria of uniform convergence, for every "\epsilon >0,\exist N\in \mathbb{N}" such that "n\geq N , m \geq N" for sufficiently large "m,n \implies"

"|S_{n}(x)-S_{m}(x)|<\epsilon , \forall x \in \mathbb{R}" "\iff" "|\sum_{r=m+1}^{n} a_r \sin(rx)|<\sum_{r=m+1}^{n}|a_r \sin(rx)|" (since, from general triangle inequality),thus

"|\sum_{r=m+1}^{n} a_r \sin(rx)|<\sum_{r=m+1}^{n}|a_r \sin(rx)|=\sum_{r=m+1}^{n}|a_r|| \sin(rx)|" .We also know that, always "|\sin(rx)|<1" "\iff \sum_{r=m+1}^{n} |a_r||\sin(rx)|<\sum_{r=m+1}^{n} |a_r|=\sum_{r=m+1}^{n} a_r" (since, "a_r >0 ,\forall r \in \mathbb{N}" .Hence choose "\epsilon = \sum_{r=m+1}^{n} a_r \iff \epsilon = \sum_{r=m+1}^{n} a_r> a_n (n-m)" ("\because \{a_n\}" is decreasing sequence),thus "a_n \rightarrow 0 ," as "n \rightarrow \infty" (since, "\epsilon" is finite positive real number).

Hence, we are done.